OpenStudy (anonymous):

solve by using the properties of rational exponents: (a^-1/3 b^1/2) (a^-1/2 b^3/5)^10

OpenStudy (campbell_st):

ok... so use the power of a power in the 2nd set of brackets \[(x^a)^b = x^{a \times b}\] so it becomes \[a^{-\frac{1}{3}}b^{\frac{1}{2}} \times a^{-\frac{1}{2} \times 10}b^{\frac{3}{5} \times 10}\] once you have that sorted you need to use the multiplication law for same base\[x^a \times x^b = x^{a + b}\] hope this helps.